OpenStudy (anonymous):

how do you solvve log sub4(x-2)- log sub4(x-3)=2

OpenStudy (anonymous):

just assume log sub4(x-2)=a and then solve it

OpenStudy (anonymous):

\[\log_{4} [(x -2)/(x -3)] = \log_{4} 16\] So, (x-2)/ (x-3) = 16 or, x-2 = 16x - 48 or, 15x = 46 so, x = 46 / 15

OpenStudy (anonymous):

howd u get 46? im confused what do you do fisrt?

OpenStudy (anonymous):

yes the answer is right

OpenStudy (anonymous):

the answer is 9/2 idk how

OpenStudy (anonymous):

9/2 is may be misprinting. 46/15 is the right answer. Place that value in your original equation (that way you can check your answer)....

OpenStudy (anonymous):

ok